# Combine and conquer: model averaging for out-of-distribution forecasting

### v2

@@ Abstract
Travel behaviour modellers have an increasingly diverse set of models at their disposal, ranging from traditional econometric structures to models from mathematical psychology and data-driven approaches from machine learning. A key question arises as to how well these different models perform in prediction, especially when considering trips of different characteristics from those used in estimation, i.e. out-of-distribution prediction, and whether better predictions can be obtained by combining insights from the different models. Across two case studies, we show that while data-driven approaches excel in predicting mode choice for trips within the distance bands used in estimation, beyond that range, the picture is fuzzy. To leverage the relative advantages of the different model families and capitalise on the notion that multiple ‘weak’ models can result in more robust models, we put forward the use of a model averaging approach that allocates weights to different model families as a function of the distance between the characteristics of the trip for which predictions are made, and those used in model estimation - in our case, we use trip distance as the sole characteristic for this. Overall, we see that the model averaging approach gives larger weight to models with stronger behavioural or econometric underpinnings the more we move outside the interval of trip distances covered in estimation. Across both case studies, we show that our model averaging approach obtains improved performance both on the estimation and test data, and crucially also when predicting mode choices for trips of distances outside the range used in estimation.
@@ 1 Introduction
Inference models and forecasting models are evaluated differently by analysts. When building an inference model, an analyst is keen to ensure that the model generates behaviourally plausible insights into explanatory factors and their relative impacts. When building a forecasting model, the focus is much more on a model’s ability to generate accurate predictions of behaviour. Importantly, in forecasting, we predict the behaviour in new settings, meaning that the models have to generalise out-of-sample. As some of the settings we wish to make predictions for may be outside the range of level-of-service values that the models have been estimated on, it is important for forecasting models to also generalise out-of-distribution.
@@
Travel behaviour analysts are increasingly interested in a more diverse set of models, covering not just econometric structures, but also models from mathematical psychology (e.g. Hancock et al., 2021, 2020a), and increasingly machine learning techniques (e.g. Alwosheel et al., 2021; Wong and Farooq, 2021; Sfeir et al., 2022; Yan et al., 2025; Wang et al., 2020). This is especially true given the growing reliance on highly granualar data on travel behaviour, for example that collected using GPS tracking (e.g. Liu et al., 2024; Tsoleridis et al., 2025).
With the typical model evaluation approach focussing on the ability of models to replicate the behaviour in the empirical data as well as possible, it should then come as no surprise that researchers in the travel behaviour field are increasingly attracted by the comparatively good prediction performance of machine learning approaches (cf. Hagenauer and Helbich, 2017).
@@
Of course, the entire purpose of building a forecasting model is not to replicate choices that have already been made, but to make predictions for new settings. Despite the growing interest in machine learning approaches, the dominant - although not necessarily evidenced-based - view is that forecasting models with a solid behavioural underpinning, such as Random Utility Maximisation (RUM) based discrete choice models, are better equipped to predict behaviour under new settings than are models with a weak or no behaviour underpinning, such as e.g. machine learning models (cf. van Cranenburgh et al., 2022).
However, what is currently less well understood is how to value and incorporate model performance, i.e. the model fit on the empirical data, in the process of developing forecasting models. What is clear is that good model performance is insufficient to establish that a model will make good out-of-distribution predictions. For that, the model must have captured underlying relationships that generalise beyond the observed data. The choice modelling literature has at times recognised that more advanced models that offer a better fit on the empirical data do not necessarily lead to better forecasts (see e.g. Fox et al., 2014). However, the question of how to develop forecasting models considering the model’s performance and behavioural underpinning is still open. On the one hand, the fact that model parameters are estimated by maximising the model performance on the empirical data, i.e. how well they replicate current choices, is somewhat at odds with the aim of forecasting, i.e. to generalise out-of-distribution. On the other hand, it is intuitive that the performance of a model on the empirical data still pertains to relevant information on its ability to generalise out-of-distribution.
The aim of this paper is twofold. First, we aim to shed light on the prediction performance of forecasting models with varying levels of behavioural underpinning as a function of the “distance” between the training data and forecasting scenarios for models. Second, we aim to develop an approach that improves prediction performance in forecasts through a model averaging approach that assigns different weights to different models depending on this “distance". In our work, we use trip distance as the key variable, but the approach works with any variable that goes out of distribution, not only geographic distance.
@@ 2.1 Generic model averaging approach
In model averaging (see e.g. Hancock et al., 2020b), which can be viewed as a form of stacking, first introduced in the seminal work of Wolpert (Wolpert, 1992), which established the conceptual basis for ensemble methods combining multiple predictive models. Specifically, we first estimate $M$ different models on the data, with model $m$ giving a likelihood $L_{n,t,m}\left(\Omega_{m}\right)$ for the choice observed by decision maker $n$ in choice situation $t$, using a vector of estimated model parameters $\Omega_{m}$. Let us further define $L_{n,m}\left(\Omega_{m}\right)$ to be the likelihood for the entire sequence of $T_{n}$ choices observed for person $n$ according to model $m$.
@@ 2.2 Model averaging for out-of-distribution prediction
Model averaging typically estimates each model on the entire data and then computes the weights by considering how well each model fits for each of the data points, estimating such weights at the person or observation level. This could, for example, include decision-maker characteristics such as age, gender and income and trip characteristics such as purpose and trip distance.
In the present paper, we focus on trip distance in particular. This is to some extent a pragmatic choice given the proof-of-concept nature of our work. It allows us to very easily split the data and define what constitutes within-distribution and out-of-distribution, with a view to demonstrating the performance of the model averaging approach.
In addition however, trip distance has a number of appealing characteristics in this context. First, distance is not typically used as an explanatory variable in the mode choice models themselves, which instead rely on the closely correlated variables such as time and cost. Second, distance can explain much of the similarity between trips, along multiple dimensions at the same time. This way, the model averaging structure can learn which model works well for what trip distance, without having to rely on multiple covariates, given that this single variable condenses other information into a single factor.
@@
In the present paper, we propose a different approach that enables us to examine the weight assigned to individual models in out-of-distribution predictions. Again, using distance as an example, let us assume that for the available estimation data, we have minimum and maximum distances $d_{min}$ and $d_{max}$, i.e. $D_{data}=\left[d_{min},d_{max}\right]$. We would then define a narrower interval $D_{train}=\left[d_{a},d_{b}\right]$, where $d_{min}<d_{a}<d_{b}<d_{max}$. In the estimation of the individual sub-models, we then use only a subset of the range of the available data, specifically those observations where $d_{n,t}\in D_{train}$. The model averaging structure on the other hand will be estimated on the entire data, i.e. still maximising e.g. Equation 2. This means that the probabilities for the individual models will be calculated for all trips in the data, but using parameters $\hat{\Omega}_{m}$ obtained from optimisation on only a part of the data (where $d_{n,t}\in D_{train}$). Model averaging learns the role of trip characteristics in determining which model should obtain more weight for a specific trip. These are grouped together in the vector $z_{n,t}$ in Equation 4. By splitting the data in the way described above, we can now define new variables to be used in Equation 4 that help characterise whether a given observation is within $D_{train}$, and if not, how far outside the training data it is located.
@@ 3.1.1 DECISIONS data
The first dataset comes from a large-scale survey conducted as part of the DECISIONS project carried out by the Choice Modelling Centre at the University of Leeds (Calastri et al., 2020). We specifically use the GPS tracking component of this survey, where the data used for this work corresponds to the observed mode choice behaviour.
After extensive data cleaning and data enrichment (Tsoleridis et al., 2022), 10,990 trips made by 415 individuals remained. For each trip, individuals travelled by one of six modes: car (48%), bus (14%), rail (5%), taxi (3%), cycling (4%) or walking (26%). The data covers a very wide range of trip distances, going from just under 100 metres to just under $106$ kilometres, with a mean distance of $8.2$ kilometres and a median distance of $4.5$ kilometres.
@@ 3.1.2 LPMC data
We use a sample of $80,943$ trips made by $31,921$ individuals. This dataset contains four alternatives: walking (18%), cycling (3%), public transport (35%, grouping together bus and rail) and driving (44%, grouping together private car and taxi rides). The data again covers a very wide range of trip distances, going from just under 100 metres to just over $40$ kilometres, with a mean distance of $4.6$ kilometres and a median distance of $2.8$ kilometres.
Attributes of the alternatives used in the models include in-vehicle travel time, out-of-vehicle travel time, interchanges, and travel cost, along with two socio-demographic variables, namely car ownership and driving license status.
@@ 3.2 Estimation and test data
In addition to these divisions, we further retain $20\%$ of any trips, independent of their distance, for out-of-sample test. The splitting of the sample into estimation, test (and additionally validation data for MLP and XGB) was done at the person/household level, so as to avoid data leakage. In other words, each person only contributed data to either the estimate data or the test data. For those individuals who contributed to the estimation data, their observations in distance segments 1 and 10 were thus not used in our analysis. The resulting division of the data is explained in Figure 1, with the resulting sample sizes shown Table 1. In practical work, analysts would be unlikely to not include all deciles (intervals of 10% width) in the MA estimation work - we made these decisions with a view to validating the model idea. Similarly, a real-world application would not set aside the test data, and thus also avoid deleting observations in the estimation data for trips that are in segments 1 and 10 (see the earlier point about data leakage).
@@ 3.3 Model types
The following individual models are used in our analysis, combining structures from choice modelling, mathematical psychology and machine learning. Many other model types could have been considered [^1] A notable omission in our list is mixture models such as mixed logit or latent class - any advantages that such models have in estimation largely/completely disappear in prediction as the random heterogeneity needs to be averaged out. The only exception comes with the use of posterior distributions, which are not applicable in the context of out-of-sample prediction. - we limited ourselves to this set with a view to demonstrating the potential benefit of our proposed approach.
@@
The NL and DFT models were coded and estimated using Apollo (cf. Hess and Palma, 2019), while the MLP and XGB models were implemented in Python. The MLP is implemented in PyTorch (Pedregosa et al., 2011) using a custom to handle the availability conditions, while the XGBoost model is implemented directly with the xgboost package (Chen and Guestrin, 2016). For both models, a hyperparameter grid search was conducted to optimise performance. We used a 5-fold Grouped cross-validation procedure to account for the panel structure of the data (see Table 2). The hyperparameter ranges for the XGB model were informed by the results of Hillel (2021).
@@
The number of hidden layers and nodes per layer influences the model’s capacity to learn complex patterns, while the batch size affects the stability and efficiency of the gradient updates. The learning rate controls how quickly the model updates its weights during training, while the L2 regularisation helps prevent overfitting by penalising large weight values. Since the best-performing hyperparameters were highly similar across both datasets, we opted to use a single configuration. The best-performing configuration found was two hidden layers with 30 nodes each, a learning rate of 0.001, a batch size of 250 and an L2 regularisation strength of 0. Because each training run of an MLP can yield slightly different outcomes, we trained the model 100 times using the best-performing configuration and averaged its predictions to account for this variability.
@@
For the XGBoost model, the hyperparameter search included the maximum tree depth, the number of estimators, the minimum loss reduction, the minimum child weight, the subsample fraction, the fraction of features considered per split, regularisation strengths (L1 and L2), and the maximum delta step. The maximum tree depth controls the complexity of each tree, with deeper trees allowing more detailed partitions at the risk of overfitting. The number of estimators determines how many boosting rounds are performed, thereby influencing the overall capacity of the model. The minimum loss reduction acts as a pruning mechanism by preventing splits that do not provide sufficient improvement to the objective function, while the minimum child weight enforces a lower bound on the sum of instance weights required in a leaf, discouraging overly specific partitions. The subsample fraction regulates the proportion of training samples used for each boosting round, and the feature subsampling parameter (colsample by tree) increases model robustness by introducing randomness at the feature level. L1 and L2 regularisation terms penalise overly large weights and help stabilise the model under strong regularisation. Finally, the maximum delta step restricts the weight update size, improving convergence stability in cases of unbalanced data. Unlike the MLPs, the optimal hyperparameters for the XGBoost model were not identical across the datasets. Both the London and Decision data shared a fixed learning rate of 0.01, a minimum loss reduction of 0.6, a minimum child weight of 20, and a subsample fraction of 0.7. However, the London configuration used a feature subsampling rate of 0.75, a maximum tree depth of 6, 1,200 boosting rounds, and regularisation strengths of $L_{1}=20$ and $L_{2}=40$, with a maximum delta step of 6. In contrast, the Decision configuration relied on full feature sampling (1.0), a deeper maximum tree depth of 9, 800 boosting rounds, and lighter $L_{1}$ regularisation of 10 (with $L_{2}=40$), combined with a smaller maximum delta step of 4. As with the MLP model, we repeated the training 100 times using the optimal configuration and averaged their predictions to account for variability in the model training. This is common practice, as when training MLPs, a substantial share gets stuck in inferior solutions.
@@
Hidden layers: $\{(10,5),(10,10),$ (20,20),(30,30)}\;(20,20),(30,30)\} Batch size: $\{250,500,1000\}$ Learning rate: $\{0.01,0.001,0.0001\}$ $L_{2}$ regularisation: $\{0,0.1,0.5,1\}$ Max. tree depth: $\{6,9\}$ Number of estimators: $\{800,1200\}$ Min. loss reduction: $\{0.2,0.6\}$ Min. child weight: $\{10,20\}$ Subsample fraction: $\{0.7,0.8\}$ Colsample by tree: $\{0.75,1.0\}$ $L_{1}$ regularisation: $\{10,20,40\}$ $L_{2}$ regularisation: $\{10,20,40\}$ Max. delta step: $\{4,6\}$

### v3

@@ Abstract
Travel behaviour modellers have an increasingly diverse set of models at their disposal, ranging from traditional econometric structures to models from mathematical psychology and data-driven approaches from machine learning. A key question arises as to how well these different models perform in prediction, especially when considering trips of different characteristics from those used in estimation, i.e. out-of-distribution prediction, and whether better predictions can be obtained by combining insights from the different models. We focus on trip distance as a key example of a variable where the application context might go beyond the estimation data. Across two case studies, we show that while data-driven approaches excel in predicting mode choice for trips within the distance bands used in estimation, beyond that range, the picture is fuzzy. To leverage the relative advantages of the different model families and capitalise on the notion that multiple ‘weak’ models can result in more robust models, we put forward the use of a model averaging approach that allocates weights to different model families as a function of the distance between the characteristics of the trip for which predictions are made, and those used in model estimation. Overall, we see that the model averaging approach gives larger weight to models with stronger behavioural or econometric underpinnings the more we move outside the interval of trip distances covered in estimation. Across both case studies, we show that our model averaging approach obtains improved performance both on the estimation and test data, and crucially also when predicting mode choices for trips of distances outside the range used in estimation.
@@ 1 Introduction
Inference models and forecasting models are evaluated differently by analysts. When building an inference model, the focus is on ensuring that the model generates behaviourally plausible insights into explanatory factors and their relative impacts. In contrast, when building a forecasting model, the emphasis is on the model’s ability to generate accurate predictions of behaviour. Importantly, forecasting involves predicting behaviour in new settings, meaning that the models have to generalise out-of-sample. This may relate, for example, to specific combinations of time and cost that were not observed in the estimation data. In addition, some prediction settings may lie outside the range of attribute values on which the models were estimated. This could involve higher costs, longer travel times, or, more intuitively, trips over longer distances [^1] We do not consider issues of spatial or temporal transferability (see e.g. Fox et al., 2014), nor the introduction of new modes or attributes. Instead, we focus on situations where attribute levels, or combinations thereof, fall outside the distribution observed in the estimation data. Our analysis is further restricted to continuous attributes, for which the impact of new attribute levels can be evaluated based on the estimated model parameters.. In such cases, it is important for forecasting models to also generalise out-of-distribution.
@@
Travel behaviour analysts are increasingly interested in a more diverse set of models, covering not just econometric structures, but also models from mathematical psychology (e.g. Hancock et al., 2021, 2020b), and increasingly machine learning techniques (e.g. Alwosheel et al., 2021; Wong and Farooq, 2021; Sfeir et al., 2022; Yan et al., 2025; Wang et al., 2020). This is especially true given the growing reliance on highly granular data on travel behaviour, for example that collected using GPS tracking (e.g. Liu et al., 2024; Tsoleridis et al., 2025). With the typical model evaluation approach focussing on the ability of models to replicate the behaviour in the empirical data as well as possible, it should then come as no surprise that researchers in the travel behaviour field are increasingly attracted by the comparatively good prediction performance of machine learning approaches (cf. Hagenauer and Helbich, 2017).
@@
Of course, the entire purpose of building a forecasting model is not to replicate choices that have already been made, but to make predictions for new settings. Despite the growing interest in machine learning approaches, the dominant, although not necessarily evidenced-based, view is that forecasting models with a solid behavioural underpinning, such as Random Utility Maximisation (RUM) based discrete choice models, are better equipped to predict behaviour under new settings than are models with a weak or no behaviour underpinning, such as e.g. machine learning models (cf. van Cranenburgh et al., 2022).
However, what is currently less well understood is how to value and incorporate model performance, i.e. the model fit on the empirical data, in the process of developing forecasting models. Performance on empirical (estimation) data is often taken as an indication of a model’s ability to generalise, despite evidence that more flexible models that achieve better in-sample fit do not necessarily lead to improved forecasts (see e.g. Fox et al., 2014; Hess and van Cranenburgh, 2025). Moreover, when behavioural patterns in out-of-distribution settings differ from those observed in the estimation data, a good in-sample fit alone may not translate into strong out-of-distribution performance. In such cases, models with a stronger behavioural underpinning may be better suited to capturing underlying decision mechanisms and could therefore generalise more robustly to new scenarios.
The contribution of this paper is twofold. First, we aim to shed light on the prediction performance of forecasting models with varying levels of behavioural underpinning as a function of the “distance” between the training data and forecasting scenarios for models. Second, we develop an approach to improve prediction performance through model averaging, where different weights are assigned to different models depending on this “distance". As highlighted above, there are many ways in which prediction settings could go outside the (multi-dimensional attribute) distribution covered by the estimation data. The focus of our work is on values of continuous attributes. We specifically use trip distance for this. It is a natural way of defining the context of a trip, and other key variables such as travel times and cost would be correlated with distance, thus also going out-of-distribution for trips that are longer or shorter than those included in the estimation data. There are also policy-relevant motivations for using distance as an example. In particular, forecasting behaviour beyond the range of observed trip distances is of interest in settings where long-distance trips are underrepresented in estimation data. While we use trip distance as the primary variable, the proposed approach is conceptually general and could be applied to other attributes that extend beyond the range observed in the estimation data. Assessing its performance in such settings is left for future work.
@@ 2.1 Generic model averaging approach
In model averaging (see e.g. Hancock et al., 2020a), which can be viewed as a form of stacking, first introduced in the seminal work of Wolpert (Wolpert, 1992), which established the conceptual basis for ensemble methods combining multiple predictive models. Specifically, we first estimate $M$ different models on the data, with model $m$ giving a likelihood $L_{n,t,m}\left(\Omega_{m}\right)$ for the choice observed by decision maker $n$ in choice situation $t$, using a vector of estimated model parameters $\Omega_{m}$. Let us further define $L_{n,m}\left(\Omega_{m}\right)$ to be the likelihood for the entire sequence of $T_{n}$ choices observed for person $n$ according to model $m$.
@@ 2.2 Model averaging for out-of-distribution prediction
Model averaging typically involves estimating each model on the full dataset and then computing weights based on how well each model fits for each of the data points, estimating such weights at the person or observation level. These weights can depend on decision-maker characteristics such as age, gender, and income, as well as trip characteristics such as purpose and trip distance.
In the present paper, as outlined in the introduction, we focus specifically on trip distance. This is partly a pragmatic choice, reflecting the proof-of-concept nature of our work. It allows for a straightforward definition of in-distribution and out-of-distribution settings through a simple data split, facilitating a clear assessment of the proposed model averaging approach. It should be noted that this definition of out-of-distribution prediction may favour models with relatively high bias and low variance, as such models often extrapolate more smoothly beyond the observed range (Bishop and Nasrabadi, 2006).
Trip distance also has several appealing characteristics in this context. First, it is typically not included directly as an explanatory variable in the mode choice models, which instead rely on closely correlated variables such as time and cost. Second, distance provides a parsimonious way of capturing variation across trips, as it is correlated with multiple underlying attributes. This allows the model averaging structure to learn how model performance varies with trip context using a single, interpretable variable, rather than relying on a higher-dimensional set of covariates.
@@
In the present paper, we propose a different approach that enables us to examine the weight assigned to individual models in out-of-distribution predictions. Again, using distance as an example, let us assume that for the available estimation data, we have minimum and maximum distances $d_{min}$ and $d_{max}$, i.e. $D_{data}=\left[d_{min},d_{max}\right]$. We would then define a narrower interval $D_{train}=\left[d_{a},d_{b}\right]$, where $d_{min}<d_{a}<d_{b}<d_{max}$. In the estimation of the individual sub-models, we then use only a subset of the range of the available data, specifically those observations where $d_{n,t}\in D_{train}$. The model averaging structure, on the other hand, will be estimated on the entire data, i.e. still maximising e.g. Equation 2. This means that the probabilities for the individual models will be calculated for all trips in the data, but using parameters $\hat{\Omega}_{m}$ obtained from optimisation on only a part of the data (where $d_{n,t}\in D_{train}$). Model averaging learns the role of trip characteristics in determining which model should obtain more weight for a specific trip. These are grouped together in the vector $z_{n,t}$ in Equation 4. By splitting the data in the way described above, we can now define new variables to be used in Equation 4 that help characterise whether a given observation is within $D_{train}$, and if not, how far outside the training data it is located.
@@ 3.1.1 DECISIONS data
The first dataset comes from a large-scale survey conducted as part of the DECISIONS project carried out by the Choice Modelling Centre at the University of Leeds (Calastri et al., 2020). We specifically use the GPS tracking component of this survey, where the data used for this work corresponds to the observed mode choice behaviour. Respondents were tracked over a period of two weeks, with most trips taking place in Yorkshire.
After extensive data cleaning and data enrichment (Tsoleridis et al., 2022), 10,990 trips made by 415 individuals remained. For each trip, individuals travelled by one of six modes: car (48%), bus (14%), rail (5%), taxi (3%), cycling (4%) or walking (26%). The data covers a wide range of trip distances, going from just under 100 metres to just under $106$ kilometres, with a mean distance of $8.2$ kilometres and a median distance of $4.5$ kilometres. It thus covers both urban trips as well as inter-city trips, for example between Leeds and York.
@@ 3.1.2 LPMC data
We use a sample of $80,943$ trips made by $31,921$ individuals. This dataset contains four alternatives: walking (18%), cycling (3%), public transport (35%, grouping together bus and rail) and driving (44%, grouping together private car and taxi rides). The data again covers a wide range of trip distances, going from just under 100 metres to just over $40$ kilometres, with a mean distance of $4.6$ kilometres and a median distance of $2.8$ kilometres. This dataset covers all of London, including short-distance trips as well as journeys from one side of London to the other.
Attributes of the alternatives used in the models include in-vehicle travel time, out-of-vehicle travel time, interchanges, and travel cost, along with two sociodemographic variables, namely car ownership and driving licence status.
@@ 3.2 Estimation and test data
As outlined above, the shortest 10% and longest 10% of trips were retained for out-of-distribution testing. In both datasets, this implies that the out-of-distribution portion corresponds to the tails of the distance distribution. The upper tail is particularly extensive, covering trips between 19.52 km and 105.9 km in the DECISIONS data, and between 11.22 km and 40.1 km in the LPMC data. This represents over 80% of the distance range in the DECISIONS data and over 70% in the LPMC data. This creates a meaningful out-of-distribution setting, with inter-city trips in the DECISIONS data and cross-London trips in the LPMC data appearing exclusively in the test set, meaning that the contexts are structurally different from the estimation data.
In addition to these divisions, we further retain $20\%$ of any trips, independent of their distance, for out-of-sample test. The splitting of the sample into estimation, test (and additionally validation data for MLP and XGB) was done at the person/household level, to avoid data leakage. In other words, each person only contributed data to either the estimate data or the test data. For those individuals who contributed to the estimation data, their observations in distance segments 1 and 10 were thus not used in our analysis. The resulting division of the data is explained in Figure 1, with the resulting sample sizes shown Table 1. In practical work, analysts would be unlikely to not include all deciles (intervals of 10% width) in the MA estimation work; we made these decisions with a view to validating the model idea. Similarly, a real-world application would not set aside the test data, and thus also avoid deleting observations in the estimation data for trips that are in segments 1 and 10 (see the earlier point about data leakage).
@@ 3.3 Model types
The following individual models are used in our analysis, combining structures from choice modelling, mathematical psychology and machine learning. Many other model types could have been considered [^2] A notable omission in our list is mixture models such as mixed logit or latent class. Any advantages that such models have in estimation largely/completely disappear in prediction as the random heterogeneity needs to be averaged out. The only exception comes with the use of posterior distributions, which are not applicable in the context of out-of-sample prediction., where we limited ourselves to this set with a view to demonstrating the potential benefit of our proposed approach.
@@
The NL and DFT models were coded and estimated using Apollo (cf. Hess and Palma, 2019), while the MLP and XGB models were implemented in Python. The MLP is implemented in PyTorch (Pedregosa et al., 2011) using a custom to handle the availability conditions, while the XGBoost model is implemented directly with the xgboost package (Chen and Guestrin, 2016).
For the two data-driven models, a hyperparameter grid search was conducted to optimise performance. We used a 5-fold Grouped cross-validation procedure to account for the panel structure of the data (see Table 2).
@@
For MLP, the number of hidden layers and nodes per layer influences the model’s capacity to learn complex patterns, while the batch size affects the stability and efficiency of the gradient updates. The learning rate controls how quickly the model updates its weights during training, while the L2 regularisation helps prevent overfitting by penalising large weight values. Since the best-performing hyperparameters were highly similar across both datasets, we opted to use a single configuration. The best-performing configuration found was two hidden layers with 30 nodes each, a learning rate of 0.001, a batch size of 250 and an L2 regularisation strength of 0.
An important point to highlight is variability in results. Each training run of an MLP can yield slightly different outcomes. An analyst then needs to make a decision on how many training runs to use, and which of the results to retain. We trained the model 100 times using the configuration described above. If our focus at this stage was purely on prediction performance, we would concentrate on the best fitting models out of these runs. However, our objective here is also to consider diversity across model instances, and capture variability arising from different local optima and promote ensemble diversity. We thus and averaged the predictions across these 100 runs to account for this variability.
@@
The hyperparameter ranges for the XGB model were informed by the results of Hillel (2021). The hyperparameter search included the maximum tree depth, the subsample fraction, the fraction of features considered per split, regularisation strengths (L1 and L2), and the maximum delta step. Furthermore, we set the learning rate to 0.1 and used the default values for all other hyperparameters. During training, early stopping was used to control the number of boosting rounds (with a maximum set to 3000). Thus, training was stopped if adding more trees did not further improve performance (on the validation set) after 10 boosting rounds. About the hyperparameter: the maximum tree depth controls the complexity of each tree, with deeper trees allowing more detailed partitions at the risk of overfitting. The subsample fraction regulates the proportion of training samples used for each boosting round, and the feature subsampling parameter (colsample by tree) increases model robustness by introducing randomness at the feature level. L1 and L2 regularisation terms penalise overly large weights and help stabilise the model under strong regularisation. Finally, the maximum delta step restricts the weight update size, improving convergence stability in cases of unbalanced data. The optimal hyperparameters for LPMC were: Max_tree_depth = 4, Subsample_fraction = 0.8, Colsample_by_tree = 0.8, $L_{1}$ reg. = 10, $L_{2}$ reg. = 1, and max_delta_set = 1. The optimal hyperparameters for DECISIONS were: Max_tree_depth = 2, Subsample_fraction = 0.4, Colsample_by_tree = 0.6, $L_{1}$ reg. = 1, $L_{2}$ reg. = 20, and max_delta_set = 2.
As with the MLP model, we repeated the training 100 times using the optimal configuration and averaged their predictions to account for variability in the model training. This is common practice, as when training MLPs, a substantial share gets stuck in inferior solutions.
@@
Hidden layers: $\{(10,5),(10,10),$ (20,20),(30,30)}\;(20,20),(30,30)\} Batch size: $\{250,500,1000\}$ Learning rate: $\{0.01,0.001,0.0001\}$ $L_{2}$ regularisation: $\{0,0.1,0.5,1\}$ Max. tree depth: $\{2,3,4,5\}$ Subsample fraction: $\{0.4,0.6,0.8\}$ Colsample by tree: $\{0.4,0.6,0.8\}$ $L_{1}$ regularisation: $\{1,5,10,20\}$ $L_{2}$ regularisation: $\{1,5,10,20\}$ Max. delta step: $\{0,1,2,4\}$